Consider the following information P(A)=0.25 P(Bc)=0.40 and P(A and B)=0.08,Then P(A|B)is 0.2


1
Expert's answer
2022-05-12T12:47:01-0400
"P(A)=0.25, P(B^C)=0.40=>P(B)=0.60"

"P(A\cap B)=0.08"

Then


"P(A|B)=\dfrac{P(A\cap B)}{P(B)}=\dfrac{0.08}{0.60}=\dfrac{2}{15}"